OpenStudy (anonymous):

equivalent equation: Jerry rented a canoo from a canoe rental company for a total of $32. The company charges $20 for the first hour and $1.50 for each additional hour. the equation below can be used to find h, the number of hours the canoe was rented. 20+(h-1)(1.5)=32.

OpenStudy (nincompoop):

okay and your question is?

OpenStudy (anonymous):

is 20(1.5)+H=32 equivelant to 20+(h-1)(1.5)=32

OpenStudy (nincompoop):

no

OpenStudy (anonymous):

got it! it's 20+(1.5h-1.5)=32

OpenStudy (anonymous):

Yes , but that's ok I figured it out thank you so much

OpenStudy (nincompoop):

let us see what we can do with your formula 20 + (h-1) (1.5) = 32 we want to look for the value of h first, we need to use distribution since we have (h-1) (1.5) 20 + 1.5h - 1.5 = 32 we want to solve for h, so we have to isolate it by subtracting 20 and then adding 1.5 on both sides of the equation (the right-hand side and left-hand side or RHS and LHS) 1.5h = 13.5 now divide both sides by 1.5 to get h